Run 10 History


Date
Notes
   
Apr. 6 We continue to run well.
            ZDC rates for 62.4 GeV/n since last Monday.

Mar. 29 The 62.4 GeV/n run is progressing along very well.
            ZDC rates for 62.4 GeV/n since last Monday.

            We had 3 instances, since starting the 62.4 GeV/n run, where
            the STAR magnet tripped off, followed by a spike up (as much as
            over 5 orders of magnitude) in the vacuum at STAR.
            This is caused by electron cloud in the STAR IR, normally suppressed
            by the presence of the solenoidal field.
            In the past a trip of the STAR magnet did not cause such a large increase
            in the vacuum, but at the lower energy we are running slightly more intense
            beams, and past magnet trips (during the 200 GeV/n run) occured at times
            when the beam was not so intense. To prevent this from occurring we have
            put the STAR magnet into the permit system. If the magnet trips again the
            beam will be aborted, preventing the buildup of an electron cloud at STAR.

            STAR vacuum showing 3 spikes when the Solenoid tripped off.
            STAR background rates showing effect of vacuum on STAR.


Mar. 22 Since Saturday evening, we have been running very well.
            ZDC rates for 62.4 GeV/n since Saturday evening.

Mar. 21 The 62.4 GeV/n run is to a good start. Machine setup went
            well, running with 111x111 bunches and 1.3e9 ions/bunch from the
            very first set of physics stores. There was a power dip yesterday
            that kept physics off for about 12 hours. Otherwise we are running
            well.
            ZDC rates for 62.4 GeV/n since Friday.
Mar. 19 62.4 GeV/n Au+Au run started physics late this afternoon.
            ZDC rates for 62.4 GeV/n.

Mar. 18 200 GeV/n Au+Au run ended this morning at 8 am.
            We began setup for the 62.4 GeV/n Au+Au run this morning.
            ZDC rates for March.
            ZDC rates for February.
            ZDC rates for January.
Mar. 15 Except or an interruption in physics due to a
            strong storm that came through hardest on
            Saturday, we have had steady running.
            On Saturday we shut down for most of the afternoon, for
            about 9 hours, since there were multiple power dips occuring.
                        ZDC rates since Thursday morning.
Mar. 11 Yesterday was a maintenance day.
            We came out of maintenance and started running physics
            by 7:30pm (last night). This can be credited to the diligence
            and hard work of the maintenance and operations teams.

            At the beginning of the week we ran into a problem in which
            Yellow beam decays increased significantly.
            It was found that the transverse coupling had changed.
            On Monday we went into a development period where we looked at
            improving the rebucketing, fixing an injection synchronization
            problem, and building a new tune/coupling replay file, by
            performing an tune/coupling feedback ramp. Things were improved
            in some ways, but we still were not running as well as the
            week before. On Tuesday we reverted some of the changes made
            on Monday, and tuned up the store working point in yellow.
                        ZDC rates since last Sunday.
Mar. 8 We have been running well the past week (as
            evident from a lack of much to say in this chronology lately.)
            For the most part, what we have been focusing on this past week
            is improving rebucketing and maximizing rates in the PHENIX
            30 cm vertex window.
            Our best stores of the run have occurred in the past 10 days.
            ZDC rates since last Wednesday.
Mar. 3 This has been a busy and productive week, since my last
            entry. Here is how things looked:
            ZDC rates since last Wednesday.

Feb. 24 The past couple days have seen very good
            running. We have fit in 8 stores with high rates.
            The instabilities at transition are under control,
            after raising the chromaticities in Yellow around
            transition time.

            There has been good progress in the stochastic cooling setup
            for both rings. The last physics store had both Yellow
            and Blue longitudinal cooling on for the entire store.

            Last night was APEX and this morning we went into maintenance.
            We come on with physics again this evening.

Feb. 22 We ran well this weekend, fitting in 12 stores
            for physics.
            ZDC rates since Friday.

            One issue we are working on is instabilities now appearing
            after transition. So we have been working on increasing octupole
            strengths at transition and trying different intensities, which
            helps understand the source of the instabilities.

Feb. 19 This has been a difficult week. Beam conditions
            have not been good. The main problem was we had high losses.
            on the ramp. This is best seen the following graph, showing
            our ramp efficiency vs fill number. Over the week this
            efficiency was dropping and bunch lengths were suffering.
            The first fill number after the maintenance was 11697.
            Ramp Efficiency since Friday, Feb. 12.

            The source of this problem turned out to be a bug in
            the LLRF, causing the transition phase jump timing to shift
            slightly every ramp.
            ZDC rates since Friday.

Feb. 16 We had a maintenance day on Friday and by Saturday
            afternoon we were running again, although at reduced
            intensities while vacuum in RHIC recovered.
            There were two areas where we opened up the vacuum to air
            to do work on the maintenance day. Both were to repair
            stochastic cooling equipment. The Yellow longitudinal
            kickers, now repaired, were reinstalled and the Blue vertical
            pickup was repaired. The vacuum in these areas is still
            recovering, limiting the beam intensity.
            The good news is the repair of the Blue vertical pickup has
            improved the transverse cooling in Blue, so that the cooling
            rate in Blue now is equal to the rate in Yellow.
            ZDC rates since Friday morning.

Feb. 12 This has been a tough past few days.
            ZDC rates since Tuesday morning.

            Tuesday morning we went off for an hour to allow PHENIX an
            access to make a repair to their EM calorimeter.
            The following store was lost to a QLI, which was caused by
            yi10-dh0 having a sudden change in its reference. The
            regulator card on that supply had to be changed. The next
            store was also lost due to a QLI. This time it appears there
            was a timing problem in the 11b-qd1 front end computer. The
            computer was reset, clearing the problem.

            Tuesday night into Wednesday morning we ran well. Wednesday
            morning a store was cut short by a QLI that occured at 5:55am.
            At the same time an ODH alarm came in, which kept us off for the
            rest of the morning. An ODH sensor had to be replaced.
            Of course, on Wednesday Long Island was hit with a major winter
            storm. As the day progressed working to get the accelerators
            running was naturally hampered by poor weather conditions.

            In the afternoon an attempt was made to put up a store, which
            failed when another QLI came in, from the yellow quadrupole
            flattop module (from yo9-qd7). The problem was due to an
            airflow fault. This finally cleared by 3pm. During this time
            the AGS main magnet power supply tripped off. This also had
            an airflow problem (bad sensor). The AGS main magnet was on
            again by 4:40pm. By 5:16pm a store was up with physics running,
            but this didn't last long, tripping off again on the yellow
            quadrupole flattop module air flow fault at 5:21pm. Again, due
            to the weather, a temporary measure was taken in which the
            airflow interlocks were bypassed (by the system engineer). This
            didn't cause any protection to be lost, since there are redundant
            temperature interlock systems on these power supplies. In the
            process of recovering another QLI, this time from yo4-qd3 which
            had gone into oscillations. This required a repair of a contactor.
            By 9:43pm we had another store up, running physics. This store ran
            until 10:56pm, at which time a vacuum interlock from sector 4
            caused the beam to be dumped. The vacuum spiked up in the area of
            the Blue stochastic cooling longitudinal kicker. In a short amount
            of time it pumped down again enough for us to operate.

            Just before midnight we got up another store, which lasted almost a
            full four hours, but was lost due to another QLI (yo5-qf2-ps had
            tripped to the Off state). After recovering a successful store
            was put up that lasted the full four hours. The next store had
            rebucketing problems and had to be aborted. That was followed by
            another successful store that lasted until the start of APEX.

            Thursday night starting at 7pm was for APEX. There were many
            problems during the APEX, causing much lost time.
            Starting at 7am Friday we went into a scheduled maintenence.

Feb. 9 We ran well over the past day. Yesterday during
            the day we were off for 1 1/2 hours when the abort
            kickers prefired (at the beginning of a store). It
            took time to recover since this caused a real quench.
            At 1pm the beam was interlocked by a faulty beam loss
            monitor.
            We are now running routinely with Blue longitudinal,
            Blue transverse, and Yellow transverse stochastic cooling.
            Cooling is not quite as effective as it could be due to
            a problem with the Blue transverse pick-up.
            Blue with longitudinal cooling.
            Yellow with no longitudinal cooling.
Feb. 8 We ran all weekend, although it was not completely
            smooth and efficient running. The time between stores
            averaged over 1.5 hours. Sunday day was a difficult day,
            in which we were struggling with instablities at
            transition. There were also some human errors made.
            Some incorrect transition timing settings were put in
            for a couple of the ramp attempts.
            We ran with both Yellow and Blue transverse stochastic
            cooling this weekend. Yellow is performing very well, while
            Blue does cool, it is not as effective due to a known problem
            with the pick-up. This is planned to be repaired on this
            Wednesday's maintenence.
            As part of the ramp development done for the STAR magnet
            polarity reversal, we moved the Yellow working point (betatron
            tune settings) to a location that recent simulations suggested
            would have more momentum aperture. This has proved to be
            a good change, since we have been able to increase the
            voltages on the 200 MHz cavities over the weekend, improving
            the rebucketing for both Blue and Yellow.
Feb. 5 Yesterday we developed a new ramp with the STAR
            magnet polarity reversed. The time to reverse the polarity
            (which involved mechanical work in the power supply) was
            a little over 4 hours. To develop the new ramp took 3
            RHIC ramps - one with tune feedback, one to do feedforward,
            and a third to fine tune things with slightly higher intensity.
            Physics was on again by 8:30pm (off a total of 12.5 hours).
            We ran 3 stores overnight.
Feb. 3 We have run 6 stores since Sunday evening.
            Early Monday morning we were held off due to a problem
            with the yi10-q89 power supply. After that fault cleared
            we got up a store that was quickly lost due to an abort
            kicker prefire. It took a few hours to recover from this
            fault. During the day Monday we shut down to allow experimental
            access. In coming on after the access we had a QLI. It took
            a few hours to recover from this fault. Yesterday we were
            off for almost five hours. A store put up at 11 am was lost
            to a QLI. This was due to a fault in the bi5-qf3 power supply.
            This morning problems with the same power supply made
            operation difficult. The power supply is being replaced.
            Today is APEX, once bi5-qf3 power supply is replaced.
Feb. 1 Physics ran all weekend.
Stores for the weekend.
Jan. 29 Physics ran all day.
            Even with the larger beta-star, we are momentum
            aperture limited in Yellow. So while we are able to
            bring on rebucketing, we cannot run at full voltage.
            The last 6 or 7 stores we have explored the level
            of voltages we can operate with.
            Polarized proton studies in the AGS continue to run behind
            RHIC stores.
Jan. 28 Yesterday was a long maintenance day.
            Physics was on again this morning at 2:40am.
            During the maintenance day the Yellow longitudinal stochastic
            cooling kickers were removed. They will be repaired in the lab
            and re-installed in a later maintenance period. The Blue transverse
            kickers were inspected and checked for leaks, but were not removed.
            They can still be used for cooling.
            Other major work was done at PHENIX. They took over 12 hours to
            make repairs to different parts of the detector.
            The AGS also took a long day to re-install the J7 stripping foil,
            used to dump the gold beams more efficiently, and to work on the
            polarimeter. Polarized proton development has begun, working
            behind RHIC stores.
Jan. 26 Physics ran all day.
            We began ramping up the rebucketing voltage today.
            As of this morning we are up to 3.3MV, working towards
            a total of 4.9MV. High high we can go may be determined by
            how much momentum aperture we find we have for this lattice.
            Studies during tonights APEX will measure this.
            APEX begins tonight at 7pm. Tomorrow is Maintenance.
Jan. 25 Physics running this weekend was cut short
            when a vacuum failure occurred on the blue transverse stochastic
            cooling kickers located in the 12 O'Clock warm section. Water
            got into the vacuum compounding the problem and causing it to
            take longer to pump down, once the leak was isolated.
            Beam went off with a QLI triggered by beam losses in Blue
            at 10:24am Sunday morning. Physics was on again at 7:27am this
            morning.
            On Friday at 4:41pm we finally got physics running on the new
            lattice with the beta-star at 0.7m (previously it was 0.6m)
            This enabled us to bring on rebucketing without large beam
            losses. Over the weekend we ramped up with beam intensity
            to be running at 1.3e9 ions/bunch with 111x111 bunches by
            Saturday evening.
            Yellow transverse stochastic cooling became operational on
            Friday evening. It has been on for every store this weekend,
            so we are gaining experience with using it.
            We have learned that without Blue transverse cooling the
            yellow beam becomes much smaller than the blue beam, causing
            luminosity to suffer later in the store. Today we will begin
            reducing the amount of Yellow cooling after 1 hour into the
            store and try to optimize the luminosity.
Jan. 22 Yesterday we spent putting up a new ramp with
            a larger beta-star. After looking over data from
            the various APEX and machine development studies
            over the past week or so, it was concluded that
            with a beta-star if 0.6m we simply didn't have
            enough momentum aperture to turn on rebucketing.
            A number of things were tried to increase the
            momentum aperture, but with limited success. So
            yesterday we put together a new ramp with a beta-star
            of 0.7m (successfully used for the dAu run) and
            overnight we worked on tuning it up.
            This morning we have put up an 84x84 bunch store with
            1e9 ions/bunch with rebucketing on (only without the
            common cavities). During the day we will work on
            putting up a 111x111 bunch store with rebucketing on
            (with common cavities? - to be discussed). This should
            put us into a good state for running physics this weekend.
Jan. 21 Yesterday was APEX, with a focus on the problems
            associated with rebucketing. Although we gained
            new insights into the problem, we did not come up
            with anything that would allow rebucketing to be brought
            on with the given optics. So it was decided later
            in the afternoon that we will develop a new ramp
            in which we go to a beta-star of 0.7m instead of 0.6m.
            The basic idea is the smaller beta-star causes the
            momentum aperture to be reduced to a level that when
            we bring on the rebucketing cavities, which causes
            the momentum spread to increase, we no longer fit
            within the existing momentum aperture.
            This work will begin this afternoon.
            In the evening we put up one store that had
            Yellow stochastic cooling (transverse) on.
            Here is how it compared to previous stores:
            PHENIX normalized coincidence rates for 4 different stores.
            The brown curve for store 11505 is with cooling turned on.
            Overnight we ran 2 more stores for physics.
Jan. 20 Here is how we ran for the last 24 hours. Stores for the last day.
            The store yesterday afternoon was longer in
            order to allow an access into the AGS. This
            was to check the AGS cold snake.
            At the end of that store, venier scans were
            done for both STAR and PHENIX.
            Early in the evening, LLRF looked at beam
            at injection to debug a problem with kicker
            timing shifts during injection. This caused
            some bunches to lose intensity during fills.
            Today, from 8am to 8pm is APEX.
Jan. 19 The weekend went very well, with a total of 16
            four hour stores for physics.
            A few hours early Sunday morning were lost due,
            initially, to an abort kicker prefire. Then there
            were problems yo1-th8, which had to be replaced.
            Here is how the week looked: Stores since Tuesday, Jan. 12
Jan. 17 Since Friday night, we have been just putting
            up stores for physics. As of 3:30 Sunday,
            we have gotten up 9 good (4 hr) stores for physics.
            When marching along, the time between stores (collisions on
            to collisions off) is 30 to 40 minutes.
            This is as fast as possible given times to ramp
            down, fill, and ramp up again.
            Here is how it looks: Stores since Friday
Jan. 15 Last night, at 7 pm, Store 11462, the stochastic
            cooling team succeeded in getting
            transverse cooling working in Yellow. Since the lattice was well
            coupled, they were able to cool both horizontal and vertical
            planes using just the vertical cooling system.
            Here is seen the cooling in action, starting at 19:00: Cooling in store 11462
            We did not recover from Wednesday's maintenance until later yesterday
            afternoon. What held us off from running again was a very
            subtle problem introduced into the LLRF during the maintenance.
            Last night and today a number of tests were
            performed in order to understand better the problem
            with rebucketing in Yellow. We continue to run
            without rebucketing because of a large beam loss
            that occurs during rebucketing in yellow.
            We will continue to work on getting rebucketing
            working next week.
Jan. 14 Yesterday was a maintenance day.
            Physics has not yet resumed, as of 8 am this morning.
            The maintenance was extended due to a problem with the
            Booster F6 extraction septum magnet. The magnet developed a short.
            This was worked on in the evening and we were able to clear it.
            The vacuum leak at the yellow longitudinal stochastic cooling
            kickers turned out to be worse than originally thought. There
            are three kickers in this unit, each individually cooled, but on
            a common system. The leaks are in the cooling lines to RHIC vacuum.
            All three kickers are leaking, suggesting a common failure mode.
            For now we are keeping the cooling lines under vacuum, preventing
            the kickers from operating.
            Another big job worked on yesterday was the J7 plunging stripping
            foil in the AGS. This is used to strip the two extra electrons from
            the Au+77 beam before being dumped into the AGS J10 dump. Although
            not required for operation, since we can dump the Au+77 beam cleanly
            without it, it does allow beam to be more effectively dumped.
            The mechanism had failed on Dec. 28, jamming in an almost inserted
            position. This did affect the beam intensity in the AGS, since we
            had to bump around it, so with it now removed we may be able to push
            intensity slightly higher (but only by 10% or so).
            As an aside, preparations are under way to begin polarized proton
            development in the AGS, behind RHIC stores. This work is development
            for the next RHIC polarized proton run (FY11?).
            This beam development will begin next week.
Jan. 13 We ran two stores for physics yesterday morning.
            An attempt to put in a third store at noon failed
            due to a LLRF problem. The afternoon was spent diagnosing
            this problem. Later in the afternoon there was a vacuum
            PLC that failed, preventing beam from being injected.
            The PLC didn't get repaired soon enough to get another store
            up before the start of APEX, originally scheduled to start
            at 8 pm. As a result of these failures APEX began at 7:14 pm.
            APEX ran all night, ending at 8 am. Today is Maintenance.
Jan. 12 We ran physics for the entire period yesterday to this
            morning. A few hours were lost due to a lead
            flow problem. In the evening a few stores were
            not successful. Two of them were lost due to beam
            losses at transition with high bunch intensities.
            The third was lost when beam debunched at injection.
Jan. 11 The basic problem causing the sextupoles to trip off was
            fixed last night. A couple short stores got put up
            earlier this morning, but they were dumped by
            a fault in Yellow 2 O'Clock TQ6.
            We are now running with 111x111 bunches with
            1.1e9 ions/bunch. But we are still running with
            the storage cavities off (no rebucketing).
            Our focus will now move to understanding why
            turning on rebucketing causes so much beam loss
            and working on commissioning stochastic cooling.
Jan. 10 The weekend has not been very good for integrating luminosity.
            Yesterday we were off most of the day due to
            a vacuum problem at the stochastic cooling kicker
            at 11 O'Clock. A store was put up at around 00:45
            this morning and another at 6:30. This store was
            lost when a sextupole tripped off. Today there has been
            a number of problems, although sextupoles tripping
            QLI's has been the most significant.
Jan. 8 The past day was another difficult struggle to get beam
            lifetimes under better control.
            Yesterday morning was given over for experimental
            access. Problems with the accesses forced a sweep
            of the experimental areas and there occured a
            problem with a switch in the access controls.
            So the 2 hour access stretched out to over 4
            hours without any beam activity.
            Before beam work could begin, a problem with the
            RHIC BPM system was enountered. This took another
            hour to sort out. Beam was not injected and ramped
            until after 4pm.
            The evening was spent working on commissioning
            a new ramp, which had modifications in it to improve
            beam lifetime at store (otherwise the ramp was no
            different than for physics stores). This work continued
            until early this morning. There remain some issues
            related to this work, most significantly rebucketing
            was left off for the most recent store.
Jan. 7 We are currently focused on improving store lifetimes, which
            are now limiting luminosity. The store put in early
            this morning was 84x84 with 1.1e9, but had lower ZDC
            rates than the previous physics store that had fewer
            bunches and lower intensity/bunch. This was all due
            to poor lifetimes.
            Overnight was for APEX. Four areas were studied: IR optics,
            orbit response matrix measurements, orbit at PHENIX,
            and IR corrections. These all went very well.
            Yesterday afternoon we were off due to two failures:
            sextupole tripped off (took time to recover) and
            the injection kickers started firing while ramping.
            The second problem was caused by a configuration error.
Jan. 6 Yesterday we went into an intensive 11 hour development period.
            This was admittedly not well planned out, but
            it proved to be highly productive. The main original
            goal of this development was to measure the chromaticity
            (first and second order) and dynamic aperture with a
            6x6 bunch ramp. What was found was chromaticities (first
            order) were far from where they should have been in a
            number of places. In particular the vertical chromaticity
            at Yellow store was large and negative, while it should
            be small and positive. So we chose to correct these
            problems.
            In addition, the LLRF group was ready to put in some
            fixes to problems encountered during commissioning.
            A major problem has been in the injection cogging.
            We have not been able to inject more than one bunch
            per AGS cycle. This leaves beam at injection for too
            long, causing significant debunching from IBS. They
            were able to fix this partially, so we can now inject
            3 out of the 4 bunches in the AGS in one cycle.
            Finally, work was done to improve the rebucketing,
            which is now working much better and with higher voltages.
            We changed the store lengths to be 4 hours starting last
            night. Previously we had set them to 6 hours, to give more
            setup time to the experiments. Now we want to ramp up
            intensity as fast as possible. Each new store will be
            incrementally higher intensity than the previous one.
Jan. 5 We continue to work on increasing intensity and run.
            Yesterday an access was made for Stochastic cooling
            to work on improving the signals from their pickups.
            We also worked on improving the rebucketing, adding
            two common cavities to achieve higher voltage.
            We lost one store last night due to a problem in the
            LLRF, causing a significant amount of beam to debunch.
            The last store put up was with 68x68 bunches with
            1e9 ions/bunch.
Jan. 4 We have started progressively raising intensity each store.
            The most recent store (11365, 1/4/10 at 6:14am) was
            with 68x68 bunches with 0.9e9 ions/bunch.
            Here is how the weekend stacked up. ZDC's and Intensity over the weekend.
Jan. 3 The beam loss that was preventing our going to higher
            intensity was cured today. Over the next few days
            we will progressively push the intensity up.
            Problems encountered today were system level problems.
            Early in the morning there was problem with injecting
            beam, associated with the bunch suppression configuration.
            Later in the morning there was a quench link interlock.
            This was followed by an independent controls problem with
            a power supply FEC. After clearing these problems a new
            fill was put up at 3 pm. This fill was with modest
            intensity, but with corrections put in late in the ramp
            cycle, the losses that were holding us from going to high
            intensity had dropped significantly.
Jan. 2 Overnight we ran with two modest stores. Each was with 56x56
            bunches with about 0.85e9 ions/bunch.
            Today, as we tried to raise intensity, either
            by increasing the bunch intensity or the number
            of bunches, we are inhibited by losses in the
            STAR triplets during the beta-squeeze.
            Efforts are now going into reducing losses
            and raising intensity.
Jan. 1 Last night a 56x56 store with 1e9 ions/bunch was put up,
            with rebucketing working.
            We are now in a mode where we will be putting up
            stores for physics, as much as possible. At the
            same time we will be increasing intensity and
            the number of bunches.
            During the days next week we will be working on
            stochastic cooling commissioning.
Dec. 31 Most of the day was spent working on commissioning the
            new LLRF for the storage system (200 MHz) RF cavities.
            The cavities were phased in and HL control setup, but
            the final step of getting beam rebucketed from the
            28 MHz RF into the 200 MHz RF needs to be done today.
            Venier scans were performed at both STAR and PHENIX
            yesterday. ZDC cross section's were measured to be the
            same as run Run7.
            At 1:30 am we got a 56x56 bunch store up with ~0.6e9
            ions/bunch, for experiment setup.
Dec. 30 Some progress was made in the last day, although a number of
            failures slowed us down. Yesterday morning we were
            kept off due to a bad UPS on the QPS in 1010A. After
            that was fixed we were kept off due to lead flow interlocks.
            Late in the afternoon we were able to run again and tested
            ramping with higher intensity (0.9 to 1.1e9 ions/bunch).
            Three such ramps were attempted before a LLRF problem
            kept us off for the rest of the evening.
            At 4 am, this morning, we got a 56x56 bunch store up
            with ~0.7e9 ions/bunch, for experiment setup.
Dec. 29 Yesterday was spent on experiment setup using a 56x56 bunch store
            and a first look at stochastic cooling commissioning.
            In the evening work continued on LLRF commissioning
            and beam development (increasing intensity).
            This morning a 61x61 bunch store was established
            for experiment setup.
Dec. 28 Last night we put in our first 56x56 bunch store, for experiments
            to begin setup for physics. Unfortunately is was aborted by
            an abort kicker prefire. Experiments have seen beam to begin
            setup, though, over the weekend.
            Over the weekend progress was made in the new LLRF setup.
            Ring to Ring synchronization was commissioned as well as
            store cogging and timing. Both rings have Landau cavities on.
            Still to be done is to setup rebucketing in the storage cavities.
            We are now ramping up to physics, focusing on store development.
Dec. 26 Work continued over the holiday to improve efficiencies on
            the ramp and at store.
            Yesterday, for the first time in Run 10, beams were put
            into collisions at both STAR and PHENIX (Merry Christmas!).
            Before we can work on physics stores, the LLRF needs to
            get ring to ring synchronization and store cogging working.
            That is being worked on today.
Dec. 24 Tuning up of the ramp efficiencies continued today.
            Overnight we have brought beam to store energies in both rings.
            Ramp efficieny in Blue is close to 100%, Yellow efficiency is poor.
            Transition crossings were worked on yesterday and fixed in both
            rings. LLRF attempted to get the Ring to Ring synchronization
            working. This needs more work, so we have no collisions yet.
            Currently we are focusing on Yellow ramp efficiency.
Dec. 23 Ramp development in both Blue and Yellow continued.
            With the focus on tuning up the transition crossing.
            With both rings running synchronously, Blue beam
            gets to store, although with only about 75% ramp
            efficiency. Most beam loss is just after transition.
            For Yellow, most of the beam is lost at transition.
            Today more work needs to be done with LLRF and ramp efficiency.
Dec. 22 Development of beam in Yellow has begun. LLRF setup
            was completed around 2 am this morning.
            Currently Yellow is being tuned to get beam to store.
            Blue has developed a problem, since we started
            ramping the two rings together. Now beam dies
            before getting to store. Since the problem is
            associated with ramping together, we are investigating
            possible causes under that condition.
Dec. 21 Newsday reported 26.3 inches of snow - at BNL - when all was said and done.
            Beam development got going last night.
            Yellow is now available for ramp development.
            Blue didn't reproduce after coming on again,
            so effort is going into working on both Blue and Yellow today.
            Schedule-wise, we are now a few days behind,
            but still pushing to have collisions by Tues. or Wed., this week.
Dec. 20 Early this morning the Blizzard of '09 finally shut us down.
            According to the National Weather Service, BNL measured a record
            24.9 inches of snowfall as of 7 am this morning.
            Beam development in RHIC continued until 3 am, when a cooling
            tower for the Booster RF went down, shutting down the Booster.
            The rest of today was spent getting roads cleared so we
            could re-start operations.
            Blue is in good shape, with beam accelerated up to store.
            Work began last night on increasing intensity in Blue.
            Today Yellow became available, so beam development will
            begin tonight in Yellow.
Dec. 18 Blue ramp development continued overnight.
            A problem was encountered the night before, in which
            the set tunes needed to be near the integer, causing
            the on-line model to crash. The actual tunes of (31.23, 31.216)
            were achieved by setting the tunes down near 31.05. It is still
            being investigated why the model is not getting them right, but
            a strong likelihood is new transfer functions that were loaded
            into the database over the summer shutdown (to match magnetic
            measurements). We reverted back to the old database yesterday
            and developed a new ramp, with all the same features. Overnight
            this ramp was developed and by this morning beam in Blue is
            up to store with >90% ramp efficiency.
            Yellow remains off due to a problem with the Q7 Quench Protection
            interface.
Dec. 17 Blue ramp development continued overnight, after being off all day
            for AGS Motor Generator repairs.
            The Blue ramp efficiency is improving, up to about 90%.
            Last night the PS group ran Yellow up to Store current levels.
            Today they need to test on the normal ramp for physics.
Dec. 16 Beam in Blue was accelerated to Store last night.
            PS group continued work on Yellow MMPS ramp development.
            Ramp development in Blue was expedited by using
            tune and coupling feedback for every ramp attempt and
            through the use of orbit feedback, which allowed correcting
            the orbit through transition in one ramp cycle.
            With a smaller beta-star, it is more difficult to correct
            the orbit during the beta-squeeze, so these systems helped
            speed up the process significantly.
            Michiko Minty put together a nice summary of the 1st 24 ramps in Blue.
            Blue ramp development summary.
Dec. 15 Ramp development for Blue began last night.
            Acceleration and the transition phase jump work well.
            LLRF setup for Blue is mostly complete, although they
            are still working on cogging.
            Yellow MMPS setup progressed today.
Dec. 14 Both Blue and Yellow are now setup at injection with RF capture.
            The MMPS for Blue is almost ready for ramp development.
            Yellow MMPS will be setup today for ramping.
            Tonight we expect to start ramp development with exp. magnets on.
Dec. 13 This afternoon we finally got beam circulating in Yellow.
            The problem turned out to be a quadrupole in the wrong polarity.
            Now we have begun to work on LLRF commissioning in Yellow.
            Also, last night, the MMPS group made progress in Blue, and
            we can now use Au101 again. We will begin working on ramp
            development in Blue tomorrow morning.
Dec. 12 Last night beam was injected into Yellow for the first time.
            Once synchro was set and the kickers timed in it became
            clear there was an obstruction in the area of 12 O'Clock.
            At most 3 turns of beam would survive.
            Today it was found the stochastic cooling pick-up at 12 O'Clock
            was not retracted. It would appear it was obstructing the beam.
            This evening, beam will be injected again and apertures around
            12 O'Clock and the Yellow cold snakes will be probed.
            LLRF commissioning for Yellow will be worked on tonight.
            For Blue, the PS group is still working on getting the Ramp
            to work. We have switched to using the Au100 ramp (which has
            the transition jumps at the same time for both rings) as oppposed
            to the Au101 ramp (which had them at different times).
            A distortion in the lattice around transition, to enable Blue
            and Yellow to jump transition at different times, was causing
            difficulties in tuning up the MMPS for that ramp. The Au100
            ramp doesn't have those distortions.
            So far they have almost gotten the current up to store levels, but
            more tuning is required.
Dec. 11 Work on Blue and Yellow MMPS continues.
            A PS failure yesterday took most of the day to fix
            (1st spare didn't work).
            Yellow could not run because of a problem in the Yellow valve box.
            This will be repaired today.
            One attempt to accelerate was made over night, but the ramp failed.
            Some beam was accelerated, though, so it was a useful first ramp.
            Blue injection is in good shape. Instrumentation is working well.
            All other systems are ready for ramp development.
Dec. 10 Worked on preparing Blue MMPS to ramp to store.
            Almost ready. Will finish today.
            LLRF Commissioning is coming along well.
            Not much beam development yesterday, but good thinking time
            (when not sleeping), so some improvements were still made
            (especially in "understanding").
            Some beam work was done last night, at injection.
            Blue injection optics are better understood.
            Instrumentation is in good shape and ready for ramps.
Dec. 9 DX training is finished. It took 3 quenches to get to Brho=840 Tm.
            LLRF Commissioning is coming along well.
            They worked all last night and are ready to work on acceleration.
            Injection and capture in blue is well tuned up.
            Orbits are in good shape and tunes and decoupling are set properly.
            Today we prepare the Blue MMPS for ramping.
            This evening we begin working on the ramp.
Dec. 8 Last night went very well. RF Capture was setup in Blue.
            So LLRF Commissioning is coming along well.
            Other work done included instrumentation setup.
            The orbit system is working, as well as the base band tune
            measurement system, beam transfer function measurements, and
            longitudinal monitors.
            Tonight we hope to do the first ramp in Blue.
Dec. 7 Blue injection and circulating beam was established last night.
            DX training was started, continues tonight.
Dec. 6 An attempt to inject beam was made, but problems with the AGS L10 cavity
            and with setup of the new RHIC LLRF prevented it.
            DX training today and tonight.
            We attempt to work on injection again in the early morning hours.
Dec. 5 No beam was injected last night.
            Beam sweeps and p.s. checkout took longer than expected.
            Tongight we work on injection and possibly RF capture.
Dec. 4 As of 6 pm, we remain on schedule.
            RHIC has been swept and secured.
            We expect to inject into Blue later tonight.
Dec. 4 Both Blue and Yellow at 4.5 K.
            X4 Storage RF cavity Damper was repaired.
Dec. 3 "mini-maintenance" this morning.
            Last night beam was re-established to the W-Dump.
            Blue is at 4.5 K, starting cooldown of Yellow.
Dec. 2 RHIC Cooldown remains on schedule. All of Blue is below 10 K.
Dec. 2 On Monday power for the Booster was shutdown (emergency work).
            A side affect was the D3 vacuum valve wouldn't open afterwards.
            (valves were closed as a normal precaution).
            It had to be opened manually.
            As of this morning, beam has not been restored
            due to other equipment failures in the Booster.
Dec. 1 Beam setup in AGS and to AtR dump is in good shape.
            Today cooldown to 4.5 K begins in Blue. 1st 9 am meeting.
Nov. 30 Put together plan for next week or so; see Plan from Dec.1 to Dec.8
Nov. 24 Working on merge setup in AGS.
Nov. 22 Beam has been extracted from the AGS to the W-dump.
Nov. 19 Today is an important day, LLRF begins switchover to the new system!
Nov. 19 Looks like the last day of the Dry Run. Most tasks are completed.
Nov. 16 completed first day of Dry Run. Cryo on schedule.